Please solve: 1)3x + 4/5 = 7/10
2)x9y6/(x3y2)3 + x5y3/(x3y)2
3)y=4x-7
y=-2x+5

Answer by Alan3354(69443)

Please solve: 1)3x + 4/5 = 7/10
Subtract 4/5, then divide by 3
-------------------------------
2)x9y6/(x3y2)3 + x5y3/(x3y)2
Not an equation, nothing to solve.
--------------------------------
3)y=4x-7
y=-2x+5
Since they both = y,
4x-7 = -2x+5
6x = 12
x = 2
etc